# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

### v2

@@ 5.1.3. Peer Review Process of HumanEvalNext
To ensure the accuracy and reliability of the initial version of HumanEvalNext, an independent reviewer verified all changes. This thorough review involved verifying the clarity and completeness of the problem docstrings, checking for consistency between the problem descriptions and the canonical solutions, and ensuring that both the solutions and test cases were correct and efficient. Where inefficiencies were identified, suggestions for optimization were provided. The review also scrutinized the test cases to ensure comprehensive coverage, identifying any gaps that could allow incorrect solutions to pass. While the initial creation of the benchmark took over 100 hours, the independent peer-review process required an additional 16 hours. As a result of this review, 1% of the problems were redesigned due to structural issues, 9% received additional test cases, and 15% underwent minor grammar or clarity improvements. All suggested changes were documented and reviewed by the original author, with every recommendation either implemented or refined further upon discussion. Since the peer-review involved modifications beyond the test suites, completions for all models were re-run. Despite these changes, the results remained largely consistent with the original, with 40% of models showing no change in their pass@1 scores, 50% showing a 1-2% absolute change, and only 10%—previously top performers—experiencing a 5% drop. This demonstrated that the peer-review process upheld the benchmark’s robustness while refining its quality.
5.2. Experimental Setup
@@
To assess the impact of the modifications applied to the benchmark, we examine the pass@1 performance of ten state-of-the-art open-source software code models using the original HumanEval alongside two enhanced variants, HumanEvalNext and EvalPlus (Liu et al., 2023a). We selected the top-performing models from the big code LLM leaderboards at the evaluation’s start: “NTQAI/Nxcode-CQ-7B-orpo”, “Qwen/CodeQwen1.5-7B”, “deepseek-ai/deepseek-coder-6.7b-instruct”, “TechxGenus/starcoder2-15b-instruct”, “ise-uiuc/Magicoder-S-DS-6.7B”, “Artigenz/Artigenz-Coder-DS-6.7B”, “HuggingFaceH4/starchat2-15b-v0.1”, “google/codegemma-7b-it”, “codeLlama/CodeLlama-13b-Instruct-hf”, and “Stabilityai/stable-code-3b”. For each task in the benchmark (164 total), the LLM is prompted using an instructional preamble asking the model to finish the implementation of the function requested in addition to providing the imports, function header, and function description with each request. We run the inference for the models on a cluster with one NVIDIA A100 80GB GPU and 32 CPU cores. During test execution, a timeout limit of 15 seconds is utilized per function call to disregard completions that are potentially looping forever or are considered overly inefficient with regard to the canonical solutions. Each test is executed in an evaluation suite using the precautions deployed by OpenAI.
5.3. Results
@@
Figure 7. Boxplot depicting the distribution of absolute drops in pass@1 score between HumanEval and the newly introduced HumanEvalNext benchmark, based on 10 LLMs (Table 22).
@@
Upon conducting the experiments outlined in Section 5.2, a key observation is an average decrease of 31.22% and a median decrease of 26.02% (both absolute percentages) in pass@$1$ results of HumanEval when contrasted with the newly introduced HumanEvalNext benchmark, using 10 different LLMs. Specific model outcomes are detailed in Table 22, with an overall depiction of the performance decreases displayed in Figure 7. While there are still significant performance declines when comparing HumanEvalPlus with the original HumanEval (11.28 mean and 7.05 median), the declines are substantially larger with HumanEvalNext than with EvalPlus. Overall, the findings highlight a marked reduction in model performance when measured against HumanEvalNext as opposed to the initial HumanEval benchmark. This pattern emphasizes the enhanced difficulty and refined evaluation precision offered by HumanEvalNext, which incorporates more resilient environments featuring type annotations and clearer instructions.
@@
Especially the resilience of models, e.g., deepseek-ai/deepseek-coder-6.7b-instruct, can be confirmed by evaluating the model on the new challenges presented in HumanEvalNext. When models also score relatively well in this benchmark, it highlights the models’ ability to adapt and perform competently under more demanding conditions, making HumanEvalNext a great benchmark to reveal the reliability of the capabilities of models. This finding also emphasizes the need for regular updates to benchmarks, as reliance on outdated benchmarks can misrepresent model capabilities over time, even when models claim not to train on the test data.
@@
Furthermore, an interesting observation emerges when analyzing model size and performance: bigger is not always better. Larger models such as TechxGenus/starcoder2-15b-instruct and HuggingFaceH4/starchat2-15b-v0.1 do not consistently outperform smaller models. This observation suggests that model size alone is not a definitive predictor of success in complex, edge-case-inclusive benchmarks like HumanEvalNext. For instance, looking at pass@1 scores, Techx- Genus/starcoder2-15b-instruct scores 77.4% on HumanEval but drops to 43.29% on HumanEvalNext, while ise-uiuc/Magicoder-S-DS-6.7B scores 76.8% on HumanEval (lower) but only drops to 53.66% on HumanEvalNext (higher), highlighting that increased model size does not necessarily equate to better performance in more challenging assessments.
@@
Lastly, ranking problems based on their difficulty using pass metrics per problem , confirms that HumanEvalNext presents a well-distributed range of complexity over the complete set of challenges. It also shows the increased difficulty of the benchmark, where even the easiest problems are not universally passed, with roughly 30% of the models still failing to solve them. Altogether, this distribution highlights HumanEvalNext’s effectiveness in providing a thorough assessment environment, making it an excellent tool for evaluating straightforward coding capabilities of LLMs in a lightweight manner - the main reason behind the popularity of the original HumanEval benchmark and its variants.
@@
Table 22. Performance Comparison on HumanEval Benchmarks (pass@1)
Model HumanEval HumanEvalPlus ($\Delta$) HumanEvalNext ($\Delta$)
stable-code-3b 30.72 25.60 (5.12-5.12) 1.83 (28.89-28.89)
CodeLlama-13b-instruct-hf 50.60 34.10 (16.50-16.50) 29.88 (20.72-20.72)
codegemma-7b-it 60.40 51.80 (8.60-8.60) 41.46 (18.94-18.94)
starchat2-15b-v0.1 73.80 71.30 (2.50-2.50) 43.29 (30.51-30.51)
Artigenz-Coder-DS-6.7B 75.60 72.60 (3.00-3.00) 53.66 (21.94-21.94)
Magicoder-S-DS-6.7B 76.80 71.30 (5.50-5.50) 53.66 (23.14-23.14)
starcoder2-15b-instruct 77.40 60.00 (17.40-17.40) 43.29 (34.11-34.11)
deepseek-coder-6.7b-instruct 80.22 71.30 (8.92-8.92) 58.54 (21.68-21.68)
CodeQwen1.5-7B 87.20 45.70 (41.50-41.50) 10.98 (76.22-76.22)
Nxcode-CQ-7B-orpo 87.23 83.50 (3.73-3.73) 51.22 (36.01-36.01)
6. Discussion
@@
In this section, we discuss the broader implications of our study; here, we attempt to tie our findins back to the RQs that guided our study. We, specifically, touch on what our results mean for researchers navigating the field, address the practical salability of our proposed framework, and outline threats to validity and opportunities for future work.
6.1. Implications
6.1.1. Finding the Right Benchmark
@@
6.1.2. Benchmarking the Benchmark
Our answer to RQ3, BenchFrame, can have an impact on how we measure progress in AI4SE. Our results show the significant impact of integrating the BenchFrame framework and highlight the value of peer-reviewed, validated benchmarks. With many state-of-the-art models tested on similar benchmarks, both industry and academia should adapt their evaluation methods to ensure robust results. The refinement of benchmarks across various AI4SE tasks will be critical for guiding future research and ensuring that these models can perform effectively. Although this study, specifically in RQ1, highlights the issue of data leakage in current benchmarks, it remains true that HumanEvalNext would also be affected by this issue. However, we argue that benchmarks should not be used indefinitely and should evolve to pose increasing challenges in alignment with model improvements.
6.2. A Case Study on the agentification of BenchFrame
A potential criticism of the framework proposed in RQ3 is the significant manual effort it demands compared to a more automated system. To address these concerns about the generalizability and manual labor intensity of BenchFrame, we developed an agentic pipeline designed to replicate and automate the benchmark improvement process with minimal human intervention. Our baseline for this effort was the HumanEvalNext dataset, which had previously been improved by manual human supervision. We aimed to determine whether an agentic approach could yield comparable improvements without the extensive human labor (i.e., to show noninfiriority).
The agent pipeline was structured into three distinct phases, which mirrors the manual process described in Section 5.1.2. In the first phase, the agent improved the problem description, which includes the function signature and the docstring. Using this refined description, the agent then generated an improved canonical solution. In the final phase, the agent created and set up the test cases to validate the solution. This process was applied across all the problems in the dataset, and our hypothesis was that this automated approach would be noninferior to the human-improved dataset.
@@ 6.2. A Case Study on the agentification of BenchFrame
Figure 8. A flow-chart indicating the pipeline used to improve MBPP.
@@
For our experiments, we used the OpenAI API  [^22] https://platform.openai.com, specifically, the o3-mini-2025-01-31 as the underlying model for the agent. We evaluated this new dataset on the same set of models, and the results (as presented in Table 23) indicated comparable (and even better) performance patterns. These findings provide adequate justification for how an agentic pipeline can effectively generalize and provide a semi-automated solution for benchmark improvement.
Table 23. Pass@1 results for HumanEval, HumanEvalNext, and HumanEvalNext-Agentic.
@@
stable-code-3b 30.72 2.44(28.28-28.28) 1.83 (28.89-28.89)
CodeLlama-13b-instruct-hf 50.60 18.90(31.70-31.70) 29.88 (20.72-20.72)
codegemma-7b-it 60.40 22.56 (37.84-37.84) 41.46 (18.94-18.94)
starchat2-15b-v0.1 73.80 26.83 (46.97-46.97) 43.29 (30.51-30.51)
Artigenz-Coder-DS-6.7B 75.60 31.70 (43.90-43.90) 53.66 (21.94-21.94)
Magicoder-S-DS-6.7B 76.80 31.70 (45.10-45.10) 53.66 (23.14-23.14)
starcoder2-15b-instruct 77.40 14.02 (63.38-63.38) 43.29 (34.11-34.11)
deepseek-coder-6.7b-instruct 80.22 35.37 (44.85-44.85) 58.54 (21.68-21.68)
CodeQwen1.5-7B 87.20 1.22 (86.00-86.00) 10.98 (76.22-76.22)
Nxcode-CQ-7B-orpo 87.23 31.70 (55.53-55.53) 51.22 (36.01-36.01)
@@
6.3. Evaluating the Generalizability of BenchFrame
An additional critique that can be raised regarding our approach is its generalizability. To tackle this issue, we adopted a strategy akin to HumanEval. Using our systematic observations of MBPP from the review, we apply BenchFrame to 100 problems selected from a pool of 500 MBPP test samples. The rationale behind this selection is twofold: firstly, it aims to assess a new task, specifically program synthesis, and secondly, it considers the benchmark’s popularity. During this process, we subject the function descriptions, implementations (including variable typing), and significantly extend test cases; more specifically, we go from a fixed 3 tests per problem to an average of 12.35, with a median of 11.5. We finally go through several rounds of peer review to ensure the quality of the improvements. We refer to the resulting benchmark as MBPPNext. The evaluation results of this dataset in comparison to the baseline are detailed in Table 24. We observe a comparable decline in performance similar to that noted for HumanEval, specifically, an average decrease of 13.4 percentage points across the methods evaluated. These findings further underscore the necessity for thorough peer review and verification of benchmarks, as discussed in previous sections of this paper.
Table 24. Pass@1 Results for MBPP and MBPPNext. Based on https://github.com/google-research/google-research/blob/master/mbpp/mbpp.jsonl accessed on the 14th of July 2025
@@ 6.3. Evaluating the Generalizability of BenchFrame
stable-code-3b 34.0 28.0 (-6%)
CodeLlama-13b-instruct-hf 40.0 31.0 (-9%)
codegemma-7b-it 43.0 33.0 (-10%)
starchat2-15b-v0.1 59.0 42.0 (-17%)
Artigenz-Coder-DS-6.7B 60.0 47.0 (-13%)
Magicoder-S-DS-6.7B 63.0 45.0 (-18%)
starcoder2-15b-instruct 44.0 36.0 (-8%)
deepseek-coder-6.7b-instruct 68.0 50.0 (-18%)
CodeQwen1.5-7B 53.0 40.0 (-13%)
Nxcode-CQ-7B-orpo 74.0 52.0 (-22%)
6.4. Future Work

### v3

@@ V-A3 Peer Review Process of HumanEvalNext
To ensure the accuracy and reliability of HumanEvalNext, an independent reviewer verified all changes. This thorough review involved verifying the clarity and completeness of the problem docstrings, checking for consistency between the problem descriptions and the canonical solutions, and ensuring that both the solutions and test cases were correct and efficient. Where inefficiencies were identified, suggestions for optimization were provided. The review also scrutinized the test cases to ensure comprehensive coverage, identifying any gaps that could allow incorrect solutions to pass. For the sake of completeness, we must note that the peer review process is exclusively concerned with evaluating the modifications made and does not encompass other elements of the procedure. While the initial creation of the benchmark took over 100 hours, the independent peer-review process required an additional 16 hours. As a result of this review, 1% of the problems were redesigned due to structural issues, 9% received additional test cases, and 15% underwent minor grammar or clarity improvements. All suggested changes were documented and reviewed by the original author, with every recommendation either implemented or refined further upon discussion. Since the peer-review involved modifications beyond the test suites, completions for all models were re-run. Despite these changes, the results remained largely consistent with the original, with 40% of models showing no change in their pass@1 scores, 50% showing a 1-2% absolute change, and only 10%—previously top performers—experiencing a 5% drop. This demonstrated that the peer-review process upheld the benchmark’s robustness while refining its quality.
V-B Experimental Setup
@@
To assess the impact of the modifications applied to the benchmark, we examine the pass@1 performance of ten state-of-the-art open-source software code models using the original HumanEval alongside two enhanced variants, HumanEvalNext and EvalPlus [10]. We selected the top-performing models from the big code LLM leaderboards at the evaluation’s start: “NTQAI/Nxcode-CQ-7B-orpo”, “Qwen/CodeQwen1.5-7B”, “deepseek-ai/deepseek-coder-6.7b-instruct”, “TechxGenus/starcoder2-15b-instruct”, “ise-uiuc/Magicoder-S-DS-6.7B”, “Artigenz/Artigenz-Coder-DS-6.7B”, “HuggingFaceH4/starchat2-15b-v0.1”, “google/codegemma-7b-it”, “codeLlama/CodeLlama-13b-Instruct-hf”, and “Stabilityai/stable-code-3b”. For each task in the benchmark (164 total), the LLM is prompted using an instructional preamble asking the model to finish the implementation of the function requested in addition to providing the imports, function header, and function description with each request. We run the inference for the models on a cluster with one NVIDIA A100 80GB GPU and 32 CPU cores. During test execution, a timeout limit of 15 seconds is utilized per function call to disregard completions that are potentially looping forever or are considered overly inefficient with regard to the canonical solutions. Each test is executed in an evaluation suite using the precautions deployed by OpenAI.
V-C Results
@@
Figure 7: Boxplot depicting the distribution of absolute drops in pass@1 score between HumanEval and the newly introduced HumanEvalNext benchmark, based on 10 LLMs (Table XXIII).
@@
Upon conducting the experiments outlined in subsection V-B, a key observation is an average decrease of 31.22% and a median decrease of 26.02% (both absolute percentages) in pass@$1$ results of HumanEval when contrasted with the newly introduced HumanEvalNext benchmark, using 10 different LLMs. Specific model outcomes are detailed in Table XXIII, with an overall depiction of the performance decreases displayed in Figure 7. While there are still significant performance declines when comparing HumanEvalPlus with the original HumanEval (11.28 mean and 7.05 median), the declines are substantially larger with HumanEvalNext than with EvalPlus. Overall, the findings highlight a marked reduction in model performance when measured against HumanEvalNext as opposed to the initial HumanEval benchmark. This pattern emphasizes the enhanced difficulty and refined evaluation precision offered by HumanEvalNext, which incorporates more resilient environments featuring type annotations and clearer instructions.
@@
Especially the resilience of models, e.g., deepseek-ai/deepseek-coder-6.7b-instruct, can be confirmed by evaluating the model on the new challenges presented in HumanEvalNext. When models also score relatively well in this benchmark, it highlights the models’ ability to adapt and perform competently under more demanding conditions, making HumanEvalNext a great benchmark to reveal the reliability of the capabilities of models. This finding also emphasizes the need for regular updates to benchmarks, as reliance on outdated benchmarks can potentially misrepresent model capabilities over time, even when models claim not to train on the test data.
@@
Furthermore, an interesting observation emerges when analyzing model size and performance: bigger is not always better. Larger models such as TechxGenus/starcoder2-15b-instruct and HuggingFaceH4/starchat2-15b-v0.1 do not consistently outperform smaller models. This observation suggests that model size alone is not a definitive predictor of success in complex, edge-case-inclusive benchmarks like HumanEvalNext. For instance, looking at pass@1 scores, Techx- Genus/starcoder2-15b-instruct scores 77.4% on HumanEval but drops to 43.29% on HumanEvalNext, while ise-uiuc/Magicoder-S-DS-6.7B scores 76.8% on HumanEval (lower) but only drops to 53.66% on HumanEvalNext (higher). This highlights that increased model size does not necessarily equate to better performance in more challenging assessments.
@@
Lastly, ranking problems based on their difficulty using pass metrics per problem, confirms that HumanEvalNext presents a well-distributed range of complexity over the complete set of challenges. It also shows the increased difficulty of the benchmark, where even the easiest problems are not universally passed, with roughly 30% of the models still failing to solve them. Altogether, this distribution highlights HumanEvalNext’s effectiveness in providing a thorough assessment environment, making it an excellent tool for evaluating straightforward coding capabilities of LLMs in a lightweight manner, the main reason behind the popularity of the original HumanEval benchmark and its variants.
@@
TABLE XXII: Performance Comparison on HumanEval Benchmarks (pass@1) - Values in parentheses are the $\Delta$ with the baseline HumanEval.
Model HumanEval HumanEval+ HumanEvalNext
stable-code 30.72 25.60 (5.12-5.12) 1.83 (28.89-28.89)
CodeLlama 50.60 34.10 (16.50-16.50) 29.88 (20.72-20.72)
codegemma 60.40 51.80 (8.60-8.60) 41.46 (18.94-18.94)
starchat2 73.80 71.30 (2.50-2.50) 43.29 (30.51-30.51)
Artigenz-Coder 75.60 72.60 (3.00-3.00) 53.66 (21.94-21.94)
Magicoder 76.80 71.30 (5.50-5.50) 53.66 (23.14-23.14)
starcoder2 77.40 60.00 (17.40-17.40) 43.29 (34.11-34.11)
deepseek-coder 80.22 71.30 (8.92-8.92) 58.54 (21.68-21.68)
CodeQwen1.5 87.20 45.70 (41.50-41.50) 10.98 (76.22-76.22)
Nxcode-CQ 87.23 83.50 (3.73-3.73) 51.22 (36.01-36.01)
TABLE XXIII: Performance Comparison on HumanEval Benchmarks (pass@1) - Values in parentheses are the $\Delta$ with the baseline HumanEval.
Model HumanEval HumanEval+ HumanEvalNext
stable-code 30.72 25.60 (5.12-5.12) 1.83 (28.89-28.89)
CodeLlama 50.60 34.10 (16.50-16.50) 29.88 (20.72-20.72)
codegemma 60.40 51.80 (8.60-8.60) 41.46 (18.94-18.94)
starchat2 73.80 71.30 (2.50-2.50) 43.29 (30.51-30.51)
Artigenz-Coder 75.60 72.60 (3.00-3.00) 53.66 (21.94-21.94)
Magicoder 76.80 71.30 (5.50-5.50) 53.66 (23.14-23.14)
starcoder2 77.40 60.00 (17.40-17.40) 43.29 (34.11-34.11)
deepseek-instruct 80.22 71.30 (8.92-8.92) 58.54 (21.68-21.68)
CodeQwen1.5 87.20 45.70 (41.50-41.50) 10.98 (76.22-76.22)
Nxcode-CQ 87.23 83.50 (3.73-3.73) 51.22 (36.01-36.01)
VI Discussion
@@
In this section, we discuss the broader implications of our study; here, we attempt to tie our findins back to the RQs that guided our study. We, specifically, touch on what our results mean for researchers navigating the field, address the practical salability of our proposed approach, and outline threats to validity and opportunities for future work.
VI-A Implications
VI-A1 Finding the Right Benchmark
@@
VI-A2 Benchmarking the Benchmark
Our answer to RQ3, BenchFrame, can have an impact on how we measure progress in AI4SE. Our results show the significant impact of integrating the BenchFrame approach and highlight the value of peer-reviewed, validated benchmarks. With many state-of-the-art models tested on similar benchmarks, both industry and academia should adapt their evaluation methods to ensure robust results. The refinement of benchmarks across various AI4SE tasks will be critical for guiding future research and ensuring that these models can perform effectively. Although this study, specifically in RQ1, highlights the issue of data leakage in current benchmarks, it remains true that HumanEvalNext would also be affected by this issue. However, we argue that benchmarks should not be used indefinitely and should evolve to pose increasing challenges in alignment with model improvements.
VI-B A Case Study on the agentification of BenchFrame
A potential criticism of the approach proposed in RQ3 is the significant manual effort it demands compared to a more automated system. To address these concerns about the generalizability and manual labor intensity of BenchFrame, we developed an agentic pipeline designed to replicate and automate the benchmark improvement process with minimal human intervention. Our baseline for this effort was the HumanEvalNext dataset, which had previously been improved by manual human supervision. We aimed to determine whether an agentic approach could yield comparable improvements without the extensive human labor (to show noninfiriority).
The agent pipeline was structured into three distinct phases, which mirrors the manual process described in subsubsection V-A2. In the first phase, the agent improved the problem description, which includes the function signature and the docstring. Using this refined description, the agent then generated an improved canonical solution. In the final phase, the agent created and set up the test cases to validate the solution. This process was applied across all the problems in the dataset, and our hypothesis was that this automated approach would be noninferior to the human-improved dataset.
@@ VI-B A Case Study on the agentification of BenchFrame
Figure 8: The flow-chart indicating the pipeline used to improve MBPP.
@@
For our experiments, we used the OpenAI API  [^22] https://platform.openai.com, specifically, the o3-mini-2025-01-31 as the underlying model for the agent. We evaluated this new dataset on the same set of models, and the results (as presented in Table XXIV) indicated comparable (and even better) performance patterns. These findings provide adequate justification for how an agentic pipeline can effectively generalize and provide a semi-automated solution for benchmark improvement.
TABLE XXIV: Pass@1 results for HumanEval, HumanEvalNext, and HumanEvalNext-Agentic.
@@
stable-code 30.72 2.44(28.28-28.28) 1.83 (28.89-28.89)
CodeLlama 50.60 18.90(31.70-31.70) 29.88 (20.72-20.72)
codegemma 60.40 22.56 (37.84-37.84) 41.46 (18.94-18.94)
starchat2 73.80 26.83 (46.97-46.97) 43.29 (30.51-30.51)
Artigenz-Coder 75.60 31.70 (43.90-43.90) 53.66 (21.94-21.94)
Magicoder 76.80 31.70 (45.10-45.10) 53.66 (23.14-23.14)
starcoder2 77.40 14.02 (63.38-63.38) 43.29 (34.11-34.11)
deepseek-coder 80.22 35.37 (44.85-44.85) 58.54 (21.68-21.68)
CodeQwen1.5 87.20 1.22 (86.00-86.00) 10.98 (76.22-76.22)
Nxcode-CQ 87.23 31.70 (55.53-55.53) 51.22 (36.01-36.01)
@@
VI-C Evaluating the Generalizability of BenchFrame
An additional critique that can be raised regarding our approach is its generalizability. To tackle this issue, we adopted a strategy akin to HumanEval. Using our systematic observations of MBPP from the review, we apply BenchFrame to 100 problems selected from a pool of 500 MBPP test samples. The rationale behind this selection is twofold: firstly, it aims to assess a new task, specifically program synthesis, and secondly, it considers the benchmark’s popularity. During this process, we subject the function descriptions, implementations (including variable typing), and significantly extend test cases; more specifically, we go from a fixed 3 tests per problem to an average of 12.35, with a median of 11.5. We finally go through several rounds of peer review to ensure the quality of the improvements. We refer to the resulting benchmark as MBPPNext. The evaluation results of this dataset in comparison to the baseline are detailed in Table XXV. We observe a comparable decline in performance similar to that noted for HumanEval, specifically, an average decrease of 13.4 percentage points across the methods evaluated. These findings further underscore the necessity for thorough peer review and verification of benchmarks, as discussed in previous sections of this paper.
TABLE XXV: Pass@1 Results for MBPP and MBPPNext. Based on https://github.com/google-research/google-research/blob/master/mbpp/mbpp.jsonl accessed on 14-07-2025
@@ VI-C Evaluating the Generalizability of BenchFrame
stable-code 34.0 28.0 (-6%)
CodeLlama 40.0 31.0 (-9%)
codegemma 43.0 33.0 (-10%)
starchat2 59.0 42.0 (-17%)
Artigenz-Coder 60.0 47.0 (-13%)
Magicoder 63.0 45.0 (-18%)
starcoder2 44.0 36.0 (-8%)
deepseek-coder 68.0 50.0 (-18%)
CodeQwen1.5 53.0 40.0 (-13%)
Nxcode-CQ 74.0 52.0 (-22%)
VI-D Future Work